# Gravitational waves from burdened primordial black holes dark matter

### v1

@@ I Introduction
But we also know that DM must already exist in the very early Universe. It cannot be something that just pops out recently. So if we want to invoke black holes as a solution of DM, they cannot be the usual kind of astrophysical black holes formed from supernovae. Instead, they must have a primordial origin. We call them primordial black holes (PBHs).
@@
If PBH is formed from collapse of overdensity region produced by inflationary perturbation, its mass is set by the scale of the Hubble horizon at the formation time Carr:2020gox . Consequently, the upper bound on the energy scale of inflation inferred from the cosmic microwave background radiation (CMB), which is around $10^{14}\ \rm GeV$, implies a lower bound on the mass of PBH: $M_{\rm PBH}>1\rm g$. According to the standard semiclassical calculation Hawking:1975vcx , PBHs in the mass range $1\rm g\lesssim M_{\rm PBH}\lesssim 10^{9}\rm g$ evaporated before the Big Bang Nucleosynthesis (BBN) and therefore are not subject to any constraint on their population. PBHs in this mass window cannot be DM (because we do see DM today), but their evaporation product could be DM. PBHs in the mass range $10^{9}\rm g\lesssim M_{\rm PBH}\lesssim 10^{17}\rm g$ evaporated between the BBN epoch and the present, so their population is constrained by multiple astrophysical data. PBHs in the mass range $10^{17}\rm g\lesssim M_{\rm PBH}\lesssim 10^{22}\rm g$ can be the entire DM; we call this the asteroid-mass window as a tyical asteroid observed in the solar system is in this mass range. They are heavy enough to not yet evaporate and they are small enough to evade gravitational lensing experiment. PBHs heavier than $10^{22}\rm g$ cannot be the entire DM as their population is constrained by lensing, accretion, and dynamical effects. Of course, one can also imagine a hybrid solution to the DM problem where PBHs together with some particle DM candidate could comprise the total observed DM abundance.
@@
However, it was recently shown that quantum memory burden effect could significantly suppress the evaporation of black hole after its half-decay time Dvali:2020wft . The half-decay time is defined as the time when the black hole has lost half of its initial mass due to evaporation and its Bekeinsgtein-Hawking entropy, which is defined as $S_{\rm BH}=4\pi GM^{2}$ in natural units Bekenstein:1973ur , reduces to a quarter of its initial value: $M\rightarrow M/2,\ S_{\rm BH}\rightarrow S_{\rm BH}/4$. The basic idea is that, because the information capacity of the black hole is reduced, some initially-gapless memory modes will acquire energy gaps due to the reduction of the number of soft gravitons. As a result, it costs much more energy to emit any further Hawking radiation and therefore creates a “burden” that prevents further decay of black hole. In particular, the evaporation rate is suppressed by a factor $S_{\rm BH}^{-k}$, where $k\geq 0$ and is a free parameter. As shown in Alexandre:2024nuo ; Thoss:2024hsr , this effect opens up a new mass window below $10^{10}\rm g$ where PBH evaporates partially until its half-decay time and then becomes stabilized to be DM. In this paper, we wish to compute the gravitational wave (GW) spectrum associated with the formation of PBHs in this new mass window.
@@
There are various physical mechanisms to form PBHs in the early Universe. The most trivial way is through the collapse of sufficiently large inflationary perturbation Carr:1974nx . Another natural way to form PBHs is through the collapse of Fermi-balls in a first-order phase transition (FOPT) Kawana:2021tde . Both of these formation scenarios could produce the corresponding GW signal that may be observed in future experiments Ghosh:2022okj ; Huang:2022him ; Xie:2023cwi . Our goal in this paper is to compute the GW spectrum associated with the formation of PBHs in the new mass window ($M_{\rm PBH}\lesssim 10^{10}\rm g$) where PBHs can be the entire DM thanks to the quantum memory burden effect. The detection of such GW signal could be an evidence that supports the existence of PBHs in this new mass window and explains the DM problem.
@@
The rest of this paper is organized as follows. In Sec.II, we show the mass and the population of PBHs formed from two different mechanisms: inflationary perturbation and collapse of Fermi ball from a FOPT. The calculations will take into account a correction due to partially evaporated black holes. In Sec.III, we show the GW spectrum corresponding to the new mass window below $10^{10}\rm g$ where PBHs can be the entire DM thanks to the memory burden effect. As we will see, GWs spectrum produced from inflation and FOPT have distinct features and, while the former peaks at the high-frequency regime, the latter peaks at the lower frequency regime that can be probed by upcoming experiments. We summarize our results in Sec.IV. The redshift business of GW spectrum in generic cosmological background is reviewed in Appendix A. The comparison between different contributions to GW spectrum from FOPT is given in Appendix B.
@@
We note that, while we were preparing our paper, we were aware of Kohri:2024qpd ; Barker:2024mpz that also study scalar-induced GW from burdened PBH DM, but they did not discuss about the GW from a FOPT scenario.
@@ II Mass and population of primordial black holes
In this section, we discuss the mass and population of PBHs in the two formation scenarios: PBHs formed from inflationary perturbation and PBHs formed from collapse of Fermi-balls in a FOPT. We assume PBHs form in the standard thermal history with the radiation-dominated phase (RD) before BBN and they have a monochromatic mass function. Unlike the standard results in the literature, we will now take into account the fact that PBH becomes stabilized after its half-life time due to the memory burden effect.
II.1 Primordial black holes formed from inflationary perturbation
@@
where $\gamma\simeq 0.2$ and $M_{\rm H,i}=\sqrt{45}m_{\rm pl}^{3}/4\pi^{3/2}g_{\rho,\rm i}^{1/2}T_{i}^{2}$ is the horizon mass at the time of formation with $m_{\rm pl}\sim 10^{19}\rm GeV$ is the Planck mass. After the PBH evaporates to half of its mass $M_{\rm PBH}=M_{\rm PBH,i}/2$, the evaporation is suppressed and the PBH becomes stabilized. In what follows, the notation $M_{\rm PBH}$ will always refer to the final mass of PBH.
PBHs with mass less than $10^{9}\rm g$ evaporate before BBN in the RD phase. Because the population of PBHs must be subdominant at early time in order to explain the correct DM abundance, their evaporation does not lead to significant entropy injection. Therefore, the population of PBHs at the half-life time is then
$\beta_{\rm PBH,half}=\frac{\rho_{\rm PBH,half}}{\rho_{\rm rad,half}}=\frac{\beta_{\rm PBH,i}}{2}\left(\frac{a_{\rm half}}{a_{i}}\right),$ (2)
where $\beta_{\rm PBH,i}$ is the initial PBH’s population and the subscript “half” refers to the quantity calculated at the half-decay time. After the half-decay time, we can reshift the PBH’s abundance to today and get
@@
We used the fact that $T\propto g_{\rm s}^{-1/3}a^{-1}$, $n_{\rm PBH,half}/s_{\rm half}\approx n_{\rm PBH}(t_{0})/s(t_{0})$ with $t_{0}$ is the present time, and $f_{\rm PBH}\equiv\Omega_{\rm PBH}/\Omega_{\rm DM}$ is the fraction of PBH in DM. PBHs can be the entire DM if $f_{\rm PBH}=1$. We also distinguish between $g_{\rho}$, which refers to the effective degrees of freedom (d.o.f.) in energy density, and $g_{\rm s}$, which refers to the effective d.o.f. in entropy density.
By using Eqs. 2 and 5 and $M_{\rm PBH,i}=2M_{\rm PBH}$, we can express the initial population of PBHs in terms of the final fraction of PBH in DM and PBH’s mass:
@@
We used the fact that, at the high temperature we consider here, $(g_{\rm s,half}/g_{\rm s,i})^{1/3}\sim O(1)$, $g_{\rho,\rm i}\sim 106.75$, and $\gamma\simeq 0.2$. We see that the quantum memory burden effect gives a correction factor $2\sqrt{2}(g_{\rm s,half}/g_{\rm s,i})^{1/3}$ to the standard result.
@@
In order to compute the GW spectrum from scalar-induced perturbation, we need an explicit form of the power spectrum of curvature perturbation $P_{\xi}(k)$. If we simply extrapolate this spectrum from the CMB scale to smaller scale, the spectrum is slightly red-tilted at small scale and therefore cannot produce PBHs. In order to have a decent chance of creating a sufficient population of PBHs, we must assume a blue-tilted power spectrum at small scale. There are many inflation models that could do this and we have no wish to invoke any specific model. In the following, we choose two simple, phenomenological forms of the power spectrum.
@@
where $k_{*}$ is the pivot scale where we have data from the CMB, $A_{\rm s}$ is the magnitude of scalar perturbation, $n_{\rm s}$ is the tilt of the spectrum, $\alpha_{\rm s}$ is the running spectral index, and $k_{\rm p}$ is the peak mode where PBHs form. We assume the running of the running is zero for simplicity. At the pivot scale $k_{*}=0.05\ \rm Mpc^{-1}$, we have $A_{\rm s}=10^{-10}\exp(3.045)$ and $n_{s}=0.965$ at $68\%$ C.L. Planck:2018jri . The allowed parameter space for the running spectral index is $\alpha_{\rm s}=0.011\pm 0.021$ at $95\%$ C.L. Cabass:2016ldu . In principle, one can consider an abrupt cutoff of the spectrum at the peak scale $k_{\rm p}$ where PBHs form. However, a smooth cutoff is more natural and we chose the form $\propto k^{-2}$ for $k\geq k_{\rm p}$ following Kohri:2018qtx . This narrowly peaked power spectrum, together with the fact that the PBH population is exponentially sensitive to the spectrum, induces an almost monochromatic mass function at $k_{p}$.
@@
The peak wavenumber that produces PBHs can be calculated as
@@
$\approx 7.31\times 10^{22}\ \rm Mpc^{-1}\left(\frac{\rm g}{M_{\rm PBH,i}}\right)^{1/2}$ (15)
$\approx 5.17\times 10^{22}\ \rm Mpc^{-1}\left(\frac{\rm g}{M_{\rm PBH}}\right)^{1/2}.$ (16)
@@
By using Eqs. 7, 8, 16, together with the power spectrum forms in Eq. 10 or Eq. 11, we can find the corresponding parameters of the spectrum that produces $f_{\rm PBH}=1$ for a given PBH’s mass. The results are shown in Table 1. We use these parameters for our computation of GW spectrum later on.
@@
$10^{3}$ 0.00889 $1.63\times 10^{21}$ 0.0128013 0.2228
$10^{6}$ 0.00954 $5.17\times 10^{19}$ 0.0146436 0.02391
$10^{9}$ 0.01029 $1.63\times 10^{18}$ 0.0169281 0.02579
@@
II.2 Primordial black holes formed from the collapse of Fermi balls during the first-order phase transition
@@
$M_{\rm PBH,i}\approx 1.4\times 10^{21}\rm g\ v_{\rm w}^{3}\left(\frac{\eta_{\chi}}{10^{-3}}\right)\left(\frac{100}{g_{*}}\right)^{1/4}\left(\frac{100\ \rm GeV}{T_{*}}\right)^{2}\left(\frac{100}{\beta/H}\right)^{3}\alpha_{*}^{1/4},$ (18)
@@
$\beta_{\rm PBH,i}^{\prime}\approx 1.4\times 10^{-15}v_{\rm w}^{-3}\left(\frac{g_{*}}{100}\right)^{1/2}\left(\frac{T_{*}}{100\ \rm GeV}\right)^{3}\left(\frac{\beta/H}{100}\right)^{3}\left(\frac{M_{\rm PBH,i}}{10^{15}\rm g}\right)^{3/2}.$ (19)
Here, $v_{\rm w}$ is the velocity of bubble wall, $\eta_{\chi}$ is the asymmetry parameter of fermion defined as $\eta_{\chi}\equiv(n_{\chi}-n_{\bar{\chi}})/s$, $T_{*}$ is the phase transition temperature with the associated effective d.o.f. $g_{*}$ at that time, $\beta/H$ is the relative time scale of phase transition and the Hubble expansion, and $\alpha_{*}$ is the ratio of vacuum energy density released to the radiation energy density of the thermal bath.
@@
Note that $\gamma_{\rm PT}$ is now not a constant but depends on phase transition’s parameters like in Eq. 18. By setting $h-0.67$ and using Eq. 6, we can again express the initial population of PBH in terms of its final abundance and mass as [^1] Note that Eq. 6 is completely general as all we said is that $M_{\rm PBH}=\gamma M_{\rm H}$. In light of Eq. 18, we can replace $\gamma\rightarrow\gamma_{\rm PT}$ with $\gamma_{\rm PT}$ being whatever in front of $M_{\rm H}$.:
@@
III Gravitational wave
In this section, we compute the GW spectrum associated with formation of PBHs from either inflationary perturbation or collapse of Fermi-balls in a FOPT. We will focus on the the new mass window below $10^{10}\rm g$ where PBHs can be the entire DM thanks to the quantum memory burden effect. As mentioned in the introduction, the suppression of PBH’s evaporation depends on the power $k$ of the Bekeinstein-Hawking entropy. The larger the power $k$, the stronger the suppression of Hawking evaporation. If $k=2$, PBHs in the mass range $10^{5}\rm g\lesssim M_{\rm PBH}\lesssim 10^{10}\rm g$ can be the entire DM. Whereas if $k=3$, the lower bound can be extended down to $\sim 10^{3}\rm g$, and so on. In order to keep the discussion general, we will not choose a specific value of $k$ but instead choose three benchmark points as an example: $M_{\rm PBH}=\{10^{3}\rm g,10^{6}\rm g,10^{9}\rm g\}$.
III.1 Scalar-induced gravitational wave from inflation
At first order in cosmological perturbation theory, the scalar, vector, and tensor modes are decoupled. They evolve independently. The vector mode decays as $\propto a^{-2}$ with $a$ being the scale factor, so we usually do not care about it. The magnitude of GW from the tensor mode depends on the tensor-to-scalar ratio, $r$. If $r$ is large, GW is large. The observational upper bound on $r$ from the CMB therefore puts a constraint on the GW produced from tensor modes.
At second order in perturbation theory, however, the scalar and tensor modes are coupled to each other. Because we know that there is scalar perturbation imprinted on the CMB, we know that there must be also the associated GW, no matter how small they are. We call this scalar-induced GW. As discussed in Sec.II.1, we must have an enhanced scalar perturbation at small scale in order to produce PBHs. This enhanced scalar perturbation will, in turn, produce the associated GW that we are about to compute [^2] It is insightful to note that, in contrast to the usual naive expectation (possibly coming from the intuition with linear physical theory), the GW produced at second order does not need to be smaller than GW produced at first order. This special feature comes from the nonlinear nature of general relativity which couples the scalar and tensor modes at higher order in perturbation theory. An enhanced scalar perturbation does not enhance GW at first order, but it can enhance GW at second order or higher..

### v2

@@ I Introduction
But we also know that DM must have already existed in the very early Universe. It cannot be something that just popped out recently. So if we want to invoke black holes as a solution of DM, they cannot be the usual kind of astrophysical black holes formed from supernovae. Instead, they must have a primordial origin. We call them primordial black holes (PBHs).
@@
If PBH is formed from collapse of overdense region produced by inflationary perturbation, its mass is set by the scale of the Hubble horizon at the formation time Carr:2020gox . Consequently, the upper bound on the energy scale of inflation inferred from the cosmic microwave background (CMB) radiation, which is around $10^{13}\ \rm GeV$, implies a lower bound on the mass of PBH: $M_{\rm PBH}>1\rm g$. According to the standard semiclassical calculation Hawking:1975vcx , PBHs in the mass range $1\rm g\lesssim M_{\rm PBH}\lesssim 10^{9}\rm g$ evaporated before the Big Bang Nucleosynthesis (BBN) and therefore are not subject to any constraint on their population. PBHs in this mass window cannot be DM (because we do see DM today), but their evaporation product could be DM. PBHs in the mass range $10^{9}\rm g\lesssim M_{\rm PBH}\lesssim 10^{17}\rm g$ evaporated between the BBN epoch and the present, so their population is constrained by multiple astrophysical data. PBHs in the mass range $10^{17}\rm g\lesssim M_{\rm PBH}\lesssim 10^{22}\rm g$ can be the entire DM; we call this the asteroid-mass window as a typical asteroid observed in the solar system is in this mass range. They are heavy enough to not yet evaporate and they are small enough to evade gravitational lensing experiment. PBHs heavier than $10^{22}\rm g$ cannot be the entire DM as their population is constrained by lensing, accretion, and dynamical effects. Of course, one can also imagine a hybrid solution to the DM problem where PBHs together with some particle DM candidate could comprise the total observed DM abundance.
@@
However, it was recently shown that quantum memory burden effect could significantly suppress the evaporation of black hole after its half-decay time Dvali:2020wft ; Dvali:2024hsb . The half-decay time is defined as the time when the black hole has lost half of its initial mass due to evaporation and its Bekeinsgtein-Hawking entropy, which is defined as $S_{\rm BH}=4\pi GM^{2}$ in natural units Bekenstein:1973ur , reduces to a quarter of its initial value: $M\rightarrow M/2,\ S_{\rm BH}\rightarrow S_{\rm BH}/4$. The basic idea is that, because the information capacity of the black hole is reduced, some initially-gapless memory modes will acquire energy gaps due to the reduction of the number of soft gravitons. As a result, it costs much more energy to emit any further Hawking radiation and therefore creates a “burden” that prevents further decay of black hole. In particular, the evaporation rate is suppressed by a factor $S_{\rm BH}^{-k}$, where $k\geq 0$ and is a free parameter. As shown in Alexandre:2024nuo ; Thoss:2024hsr , this effect opens up a new mass window below $10^{10}\rm g$ where PBH evaporates partially until its half-decay time and then becomes stabilized to be DM.
If this novel quantum effect of black hole turns out to be true, it would be natural and interesting to study various resulting phenomenologies. For example, DM production or baryogenesis from evaporation of burdened PBHs has been studied in Haque:2024eyh ; Barman:2024iht . A PBH dominated era can also produce gravitational wave (GW) due to number density fluctuation when the PBHs evaporate Papanikolaou:2020qtd ; Papanikolaou:2022chm , which is useful in probing primordial non-Gaussianity Papanikolaou:2024kjb ; He:2024luf , and so modified Hawking evaporation would also lead to modified GW spectrum Balaji:2024hpu ; Bhaumik:2024qzd .
@@
In this paper, we wish to compute the GW spectrum associated with the formation of PBHs in the new mass window below $10^{10}\rm g$ where PBHs can be the entire DM thanks to the quantum memory burden effect. There are various physical mechanisms to form PBHs in the early Universe. The most trivial way is through the collapse of sufficiently large inflationary perturbation Carr:1974nx . Another natural way to form PBHs is through the collapse of Fermi-balls in a first-order phase transition (FOPT) Kawana:2021tde . Both of these formation scenarios could produce the corresponding GW signal that may be observed in future experiments Ghosh:2022okj ; Huang:2022him ; Xie:2023cwi  [^1] These two PBH formation scenarios are just a few among many other possible mechanisms such as PBHs formed from quark confinement Dvali:2021byy , PBHs formation in the context of loop quantum cosmology Papanikolaou:2023crz or F(R) gravity Banerjee:2022xft , and many others. Nevertheless, the two formation scenarios from inflation and FOPT that we consider here are usually recognized as the most popular and well-motivated ones.. The detection of such GW signal could be an evidence that supports the existence of PBHs in this new mass window and explains the DM problem [^2] Other complimentary studies such as neutrinos or cosmic rays emission from burdened PBHs can also be useful to conclusively confirm the existence of such PBHs Chianese:2024rsn ; Zantedeschi:2024ram ..
@@
The rest of this paper is organized as follows. In Sec.II, we show the mass and the population of PBHs formed from two different mechanisms: inflationary perturbation and collapse of Fermi ball from a FOPT. The calculations will take into account a correction due to partially evaporated black holes. In Sec.III, we compute the GW spectrum corresponding to the new mass window below $10^{10}\rm g$ where PBHs can be the entire DM thanks to the memory burden effect. As we will see, GWs spectrum produced from inflation and FOPT have distinct features and, while the former peaks at the high-frequency regime, the latter peaks at the lower frequency regime that can be probed by upcoming experiments. We summarize our results in Sec.IV. The redshift business of GW spectrum in generic cosmological background is reviewed in Appendix A. Comparison between different contributions to the total GW spectrum from FOPT is given in Appendix B.
@@
We note that, while we were preparing our paper, we were aware of Kohri:2024qpd ; Barker:2024mpz ; Jiang:2024aju that also study scalar-induced GW from burdened PBH DM [^3] See also Franciolini:2023osw for GW associated with the formation of microscopic DM relics, including the Hawking remnants after evaporation., but they did not discuss about the GW from a FOPT scenario.
@@ II Mass and population of primordial black holes
In this section, we discuss the mass and population of PBHs in the two formation scenarios: PBHs formed from inflationary perturbation and PBHs formed from collapse of Fermi-balls in a FOPT. We assume PBHs form in the standard thermal history with the radiation-dominated (RD) phase before BBN and they have an almost monochromatic mass function. Unlike the standard results in the literature, we will now take into account the fact that PBH becomes stabilized after its half-decay time due to the memory burden effect.
II.1 PBHs formed from inflationary perturbation
@@
where $\gamma\simeq 0.2$, $M_{\rm H,i}=\sqrt{45}m_{\rm pl}^{3}/4\pi^{3/2}g_{\rho,\rm i}^{1/2}T_{\rm i}^{2}$ is the horizon mass at the formation time with $m_{\rm pl}\sim 10^{19}\ \rm GeV$ is the Planck mass and $g_{\rho,\rm i}$ is the effective relativistic degrees of freedom (d.o.f.) at temperature $T_{\rm i}$. This can also be written in a normalized form as
$\left(\frac{M_{\rm PBH,i}}{\rm g}\right)\approx 1.89\times 10^{31}\left(\frac{\gamma}{0.2}\right)\left(\frac{106.75}{g_{\rho,\rm i}}\right)^{1/2}\left(\frac{\rm GeV}{T_{\rm i}}\right)^{2}.$ (2)
After the PBH evaporates to half of its initial mass $M_{\rm PBH}=M_{\rm PBH,i}/2$, the evaporation is suppressed and the PBH becomes stabilized. In what follows, the notation $M_{\rm PBH}$ will always refer to the final mass of PBH.
PBHs with mass less than $10^{9}\rm g$ evaporate before BBN in the RD phase. Because the population of PBHs must be subdominant at early time in order to explain the correct DM abundance, their partial evaporation does not lead to significant entropy injection. Therefore, the population of PBHs at the half-decay time is
$\beta_{\rm PBH,half}=\frac{\rho_{\rm PBH,half}}{\rho_{\rm rad,half}}=\frac{\beta_{\rm PBH,i}}{2}\left(\frac{a_{\rm half}}{a_{\rm i}}\right),$ (3)
where $\beta_{\rm PBH,i}$ is the initial PBH’s population. After the half-decay time, we can reshift the PBH’s abundance to today and get
@@
We used the fact that $\rho_{\rm rad}=3sT/4$, $T\propto g_{\rm s}^{-1/3}a^{-1}$, $n_{\rm PBH,half}/s_{\rm half}=n_{\rm PBH}(t_{0})/s(t_{0})$ with $t_{0}$ is the present time, and $f_{\rm PBH}\equiv\Omega_{\rm PBH}/\Omega_{\rm DM}$ is the current fraction of PBH in DM. PBHs can be the entire DM if $f_{\rm PBH}=1$. For the numerical factors, we used $m_{\rm pl}=1.22\times 10^{19}\ \rm GeV=2.18\times 10^{-5}\rm g$, $s(T_{0})=2\pi^{2}g_{\rm s,0}T_{0}^{3}/45$ with $g_{\rm s,0}=3.94$ and $T_{0}=2.3\times 10^{-13}\ \rm GeV$, $\Omega_{\rm DM}=0.27$, $H_{0}=67\ \rm km/s/Mpc=1.43\times 10^{-42}\ \rm GeV$ Planck:2018vyg . We also distinguished between $g_{\rho}$, which refers to the effective relativistic d.o.f. in energy density, and $g_{\rm s}$, which refers to the effective relativistic d.o.f. in entropy density.
By using Eqs. 3 and 6, and $M_{\rm PBH,i}=2M_{\rm PBH}$, we can express the initial population of PBHs in terms of the final fraction of PBH in DM and PBH’s mass:
@@
We used the fact that, at the high temperature we consider here, $(g_{\rm s,half}/g_{\rm s,i})^{1/3}\sim O(1)$, $g_{\rho,\rm i}\sim 106.75$, and $\gamma\simeq 0.2$. We see that the quantum memory burden effect gives a correction factor $2\sqrt{2}(g_{\rm s,half}/g_{\rm s,i})^{1/3}$ to the standard result. This implies that, for a given final mass and population of PBHs, the initial population must be slightly greater than the usual value to account for the partial evaporation of PBHs.
@@
In order to compute the GW spectrum from scalar-induced perturbation, we need an explicit form of the power spectrum of curvature perturbation $P_{\xi}(k)$. If we simply extrapolate this spectrum from the CMB scale to smaller scale, the spectrum is slightly red-tilted at small scale and therefore cannot produce PBHs. In order to have a decent chance of creating a sufficient population of PBHs, we must assume a blue-tilted power spectrum at small scale. There are many inflation models that could do this and we have no wish to invoke any specific model. Instead, we will choose two simple, phenomenological forms of the power spectrum.
@@
where $k_{*}$ is the pivot scale where we have data from the CMB, $A_{\rm s}$ is the magnitude of scalar perturbation, $n_{\rm s}$ is the tilt of the spectrum, $\alpha_{\rm s}$ is the running spectral index, and $k_{\rm p}$ is the peak mode where PBHs form. We assume the running of the running is zero for simplicity. At the pivot scale $k_{*}=0.05\ \rm Mpc^{-1}$, we have $A_{\rm s}=10^{-10}\exp(3.045)$ and $n_{s}=0.9625$ at $68\%$ C.L. Planck:2018jri . The allowed parameter space for the running spectral index is $\alpha_{\rm s}=0.011\pm 0.021$ at $95\%$ C.L. Cabass:2016ldu . In principle, one can consider an abrupt cutoff of the spectrum at the peak scale $k_{\rm p}$ where PBHs form. However, a smooth cutoff is more natural and we chose the form $\propto k^{-2}$ for $k\geq k_{\rm p}$ following Kohri:2018qtx . This narrowly peaked power spectrum, together with the fact that the PBH population is exponentially sensitive to the spectrum, induces an almost monochromatic mass function at $k_{p}$.
@@
The peak wavenumber corresponding to the formation of PBHs can be calculated as
@@
$\approx 7.3\times 10^{22}\ \rm Mpc^{-1}\left(\frac{\rm g}{M_{\rm PBH,i}}\right)^{1/2}$ (16)
$\approx 5.16\times 10^{22}\ \rm Mpc^{-1}\left(\frac{\rm g}{M_{\rm PBH}}\right)^{1/2}.$ (17)
@@
By using Eqs. 8, 9, 17, together with the power spectrum forms in Eq. 11 or Eq. 12, we can find the corresponding parameters of the spectrum that produces $f_{\rm PBH}=1$ for a given PBH’s mass. The results are shown in Table 1. We use these parameters for our computation of GW spectrum later on.
@@
$10^{3}$ 0.00889 $1.63\times 10^{21}$ 0.0128 0.02228
$10^{6}$ 0.00954 $5.16\times 10^{19}$ 0.0146 0.02391
$10^{9}$ 0.01029 $1.63\times 10^{18}$ 0.0169 0.02579
@@
II.2 PBHs formed from the collapse of Fermi balls during a FOPT
@@
$M_{\rm PBH,i}\approx 1.4\times 10^{21}{\rm g}\ v_{w}^{3}\left(\frac{\eta_{\chi}}{10^{-3}}\right)\left(\frac{100}{g_{*}}\right)^{1/4}\left(\frac{100\ \rm GeV}{T_{*}}\right)^{2}\left(\frac{100}{\beta/H_{*}}\right)^{3}\alpha_{*}^{1/4},$ (19)
@@
$\beta_{\rm PBH,i}^{\prime}\approx 1.4\times 10^{-15}v_{w}^{-3}\left(\frac{g_{*}}{100}\right)^{1/2}\left(\frac{T_{*}}{100\ \rm GeV}\right)^{3}\left(\frac{\beta/H_{*}}{100}\right)^{3}\left(\frac{M_{\rm PBH,i}}{10^{15}\rm g}\right)^{3/2}.$ (20)
Here, $v_{w}$ is the velocity of bubble wall, $\eta_{\chi}$ is the asymmetry parameter of fermion defined as $\eta_{\chi}\equiv(n_{\chi}-n_{\bar{\chi}})/s$, $T_{*}$ is the phase transition temperature with the associated effective d.o.f. $g_{*}$ at that time, $\beta/H_{*}$ is the relative time scale of phase transition to the Hubble expansion, and $\alpha_{*}$ is the ratio of vacuum energy density released to the radiation energy density of the thermal bath.
@@
Note that $\gamma_{\rm PT}$ is now not a constant but depends on phase transition’s parameters like in Eq. 19. $h$ is the normalized Hubble constant and we can set it to be $h=0.67$ Planck:2018vyg . By using Eq. 7, we can again express the initial population of PBH in terms of its final abundance and mass as [^4] Note that Eq. 7 is completely general as all we said is that $M_{\rm PBH}=\gamma M_{\rm H}$. In light of Eq. 19, we can replace $\gamma\rightarrow\gamma_{\rm PT}$ with $\gamma_{\rm PT}$ being whatever in front of $M_{\rm H}$.:
@@
III Gravitational waves
In this section, we compute the GW spectrum associated with the formation of PBHs from either inflationary perturbation or collapse of Fermi-balls in a FOPT. We will focus on the the new mass window below $10^{10}\rm g$ where PBHs can be the entire DM thanks to the quantum memory burden effect. As mentioned in the introduction, the suppression of PBH’s evaporation depends on the power $k$ of the Bekeinstein-Hawking entropy. The larger the power $k$, the stronger the suppression of Hawking evaporation. If $k=2$, PBHs in the mass range $10^{5}\rm g\lesssim M_{\rm PBH}\lesssim 10^{10}\rm g$ can be the entire DM. Whereas if $k=3$, the lower bound can be extended down to $\sim 10^{3}\rm g$, and so on. In order to keep the discussion general, we will not choose a specific value of $k$ but instead choose three benchmark points as an example: $M_{\rm PBH}=\{10^{3}\rm g,10^{6}\rm g,10^{9}\rm g\}$.
III.1 Scalar-induced GW from inflation
At first order in cosmological perturbation theory, the scalar, vector, and tensor modes are decoupled. They evolve independently. The vector mode decays as $\propto a^{-2}$ with $a$ being the scale factor, so we usually do not care about it. The magnitude of GW from the tensor mode depends on the tensor-to-scalar ratio, $r$. If $r$ is large, GW is large. The observational upper bound on $r$ from the CMB therefore puts a constraint on the GW produced from tensor modes Maggiore:2018sht .
At second order in perturbation theory, however, the scalar and tensor modes are coupled to each other. Because we know that there is scalar perturbation imprinted on the CMB, we know that there must be also the associated GW, no matter how small they are. We call this scalar-induced GW (see Domenech:2021ztg ; Yuan:2021qgz for a review). As discussed in Sec.II.1, we must have an enhanced scalar perturbation at small scale in order to produce PBHs. This enhanced scalar perturbation will, in turn, produce the associated GW that we are about to compute [^5] It is insightful to note that, in contrast to the usual naive expectation (possibly coming from the intuition with linear physical theory), the GW produced at second order does not need to be smaller than GW produced at first order. This special feature comes from the nonlinear nature of general relativity which couples the scalar and tensor modes at higher order in perturbation theory. An enhanced scalar perturbation does not enhance GW at first order, but it can enhance GW at second order or higher..